25 kg of water at 95°C mix with 35 kg of water at 35°C, the pressure being taken as constant and temperature of surrounding being 15°C. Total available energy (before mixing) is (Cp of water= 4.2 KJ/kg.k)
a) 1085 KJ
b) 400 KJ
c) 985 KJ
d) 685 KJ
1
Expert's answer
2019-09-13T10:17:21-0400

Solution:


The available energy of 25 kg of water at 95°C:


"(A.E.)_{25} = mc_p\int_{T_0}^{T} (1-\frac{T_0}{T})dT ="


"=25\cdot{4.2}\int_{273+15}^{273+95} (1-\frac{288}{T})dT ="


"= 105[(368-288)-288\cdot{ln\frac{368}{288}}] = 987.49" kJ.


The available energy of 35 kg of water at 35°C:


"(A.E.)_{35} = mc_p\int_{T_0}^{T} (1-\frac{T_0}{T})dT ="


"=35\cdot{4.2}\int_{273+15}^{273+35} (1-\frac{288}{T})dT ="


"= 147[(308-288)-288\cdot{ln\frac{308}{288}}] = 97.59" kJ.


Total available energy (before mixing) is:


"(A.E.)_{total} = (A.E.)_{25} + (A.E.)_{35} = 987.49+97.59 = 1085" kJ.


Answer:


Total available energy (before mixing) is 1085 kJ.